# HEROS: Hierarchical Exploration with Online Subregion Updating for 3D Environment Coverage

### v2

@@ Abstract
We present an autonomous exploration system for efficient coverage of unknown environments. First, a rapid environment preprocessing method is introduced to provide environmental information for subsequent exploration planning. Then, the whole exploration space is divided into multiple subregion cells, each with varying levels of detail. The subregion cells are capable of decomposition and updating online, effectively characterizing dynamic unknown regions with variable resolution. Finally, the hierarchical planning strategy treats subregions as basic planning units and computes an efficient global coverage path. Guided by the global path, the local path that sequentially visits the viewpoint set is refined to provide an executable path for the robot. This hierarchical planning from coarse to fine steps reduces the complexity of the planning scheme while improving exploration efficiency. The proposed method is compared with state-of-art methods in benchmark environments. Our approach demonstrates superior efficiency in completing exploration while using lower computational resources.
Index Terms:
Search and rescue robots, mapping, autonomous agents.
@@ I Introduction
Autonomous exploration using ground mobile robots has received considerable attention in recent years for its wide applications, including searching and rescuing, environmental monitoring, and industrial automation. The ability of robots to navigate and explore unknown environments without human intervention enhances operational efficiency and ensures personnel safety in hazardous conditions. Although several autonomous exploration methods have been proposed over the decades [1], [2], [3], the task remains challenging in large-scale and topologically complex scenarios. In such scenarios, the problem becomes computationally complex, and the robot needs to navigate between areas in a limited time while avoiding redundant exploration.
The mainstream exploration frameworks mainly include frontier-based and sampling-based methods. In autonomous exploration, the frontier is defined as the boundary between free and unknown areas [1]. These two approaches guide the robot to cover the environment by continuously searching for frontiers in the map or randomly sampling viewpoints and evaluating their potential revenue. However, several imitations still hinder the development of autonomous exploration:
@@
Figure 1: The real-world experiment conducted in an underground garage using real ground platforms. The yellow dot denotes the start position, and the white line indicates the robot’s trajectory.
Taking inspiration from the aforementioned concerns, we propose an efficient hierarchical exploration planner to enable robots to explore complex unknown environments autonomously. First, we design a rapid environmental preprocessing method that incrementally extracts frontiers, continuously constructs the sparse graph of the environment, and samples high-quality viewpoints. To effectively represent the continuously changing unknown areas during the exploration process, we adopt a subregion-based regional division approach. This approach decomposes the entire unknown environment into several subregion cells with a hierarchical structure, which can be further subdivided to adapt to the changing unknown areas. Additionally, the Subregion Information Structure (SIS) is introduced to record basic information within the subregion space, facilitating subsequent exploration planning. Finally, the subregion is adopted as the basic planning unit, and the hierarchical planning strategy is used for efficient exploration. The hierarchical planning includes global coverage planning and local planning. The global coverage path is the shortest travel path that enters each subregion in sequence, while the local path is the viewpoint access path refined under the guidance of the global path. The above process is continuously executed until the entire environment is fully explored.
@@
  • A subregion-based regional division method with a hierarchical structure is proposed, allowing for adaptive decomposition of the subregion to capture the dynamic changes of unknown regions.
@@ II Related Works
Several methods [13], [14], [15], [16] are proposed for the concepts of regional division and hierarchical planning. However, these methods are either designed for task allocation in multi-robot collaborative exploration or utilize fixed-sized regions that cannot adapt to dynamic changes in unknown areas. In [15], the global path is introduced to provide a more comprehensive perspective for exploration. However, the greedy strategy is still adopted in the local planning stage, which can still lead to short-sighted behavior and thus reduce exploration efficiency. The concept of region hierarchical division in this paper is inspired by [14], which utilizes hgrid space for muti-UAV interaction and task assignment. Our approach of the regional division is aimed at serving the global coverage path planning, which is part of the hierarchical planning.
Learning-based techniques have also emerged as promising solutions for autonomous exploration [17], [18], [19]. These methods leverage data-driven models to predict optimal exploration strategies but often lack robustness and adaptability to unfamiliar scenarios. This leads to performance deterioration in complex and unfamiliar environments.
@@ IV-A Mapping and Frontier Detection
As in state-of-art works [10, 21], an incremental frontier detection method is employed in this work to avoid processing the entire map. Specifically, a set $\mathcal{C}_{v}$ of all voxels for which the occupation value has changed is obtained from the UFOMap, and the state of each voxel $V\in\mathcal{C}_{v}$ is checked. If a voxel $V$ is in the free state (denoted as $V_{free}$) and an adjacent voxel is in the unknown state (denoted as $V_{un}$) at the same height, then $V$ can be considered as a frontier voxel. During the exploration process, each frontier voxel located in the map updated region $\mathcal{B}_{u}$ is checked to determine if it can still be a frontier, and those invalid frontiers are removed. This method only requires checking a small number of voxels in the map, significantly reducing the workload for frontier extraction.
@@ IV-B Dynamic Sparse Graph
Given that our exploration method requires extensive path search operations, which are time-consuming in the 3D voxel map. Thus, a dynamic sparse graph $\mathcal{G}=(N,E)$ is maintained as a sparse representation of the environment. Here, $N$ represents the nodes in the graph, and $E$ denotes the undirected weighted edge connecting adjacent nodes. During the exploration, candidate nodes $N_{c}$ are generated through the random biased sampling method based on the terrain map $\mathcal{A}$. The sampled node that satisfies the conditions in Eq. 1 will undergo collision checking, and then be added to $\mathcal{G}$. The weight of the edge $E$ is set as the Euclidean distance between two nodes. All the nodes are stored in R*-tree [22] to speed up the computational efficiency. It is also important to perform collision checks for all edges within the boundary $\mathcal{B}_{u}$ during the exploration, and remove those that become invalid due to dynamic changes in the environment.
@@
where ${Accessible}(\mathcal{A})$ denotes the accessible region in $\mathcal{A}$, since all nodes should be reachable by the robot. The parameters $\delta$ and $\gamma$ represent the minimum and maximum allowable distances between $N_{c}$ and the nearest node $N_{\text{nearest}}$ in $\mathcal{G}$, respectively. This can prevent the graph from becoming over dense or sparse.
@@ IV-C Viewpoint Generation
where $\mathbf{p}_{o}$ is the position of the nearest occupied voxel, and $r_{\text{sensor}}$ is the maximum scanning radius of the LiDAR sensor. Due to the absence of obstacles within the sphere and its convex nature, any frontier voxel in the sphere can be observed from a viewpoint inside. Thus, performing expensive ray tracing on every frontier voxel to determine its visibility from the viewpoint is unnecessary. The overall generation procedure of the collision-free sphere is similar to [24]. Then, the node $N$ located inside the sphere and closest to its center is selected as the viewpoint. This search process is achieved by K-Nearest Neighbor search in $\mathcal{G}$. Fig. 3 shows the dynamic sparse graph and the viewpoints generated based on collision-free spheres.
@@ V Subregion-based Regional Division
Instead of directly solving the visit order of viewpoints or frontiers, we use subregion cells as basic planning units for global coverage. Each candidate cell, with varying levels of detail, is managed through online decomposition and updating for a multi-resolution representation of the unknown space.
Figure 4: Schematic representation of the hierarchy of subregion cells. Subregion cells can be further subdivided to represent unknown regions using smaller resolutions up to the maximum hierarchy $l_{\text{max}}$.
@@ V-B Subregion Information Structure
The Subregion Information Structure $\mathbf{SI_{i}}$ is composed of elemental information contained in each subregion cell $S_{i}$. The hierarchy level $l_{i}$ and axis-aligned bounding box $\mathbf{B}_{i}$ are recorded when initializing $S_{i}$. The proportion of known voxels $\mathbf{R}_{\text{known},i}$ and the viewpoint average position $\mathbf{C}_{\text{avg},i}$ of $S_{i}$ are also computed to locate the area to be explored. To serve subsequent global coverage path planning, the traveling cost $\mathbf{L}_{\text{cost},i}$ between $S_{i}$ and other subregions is also computed. The viewpoints $\mathbf{VP}_{i}$ within $S_{i}$ are stored in $\mathbf{SI_{i}}$ for subsequent local path optimization. In addition, each subregion cell holds a status $stat_{i}$ between “active” and “inactive”. Subregion cells containing at least one viewpoint are categorized as active, while those without viewpoints are considered inactive. Only subregions with active status are eligible for inclusion in the global coverage path planning process. Table. I presents the meaning of each data information in SIS.
@@
$stat_{i}$ ”active” or ”inactive” status of $S_{i}$
@@ V-C Online Subregion Updating
We maintain a set $\mathcal{V}$ that stores all divided subregions. At the start of the exploration, the entire unknown space $\mathcal{O}$ is divided into $S_{o}=S_{o,x}\times S_{o,y}$ cells at level $l=1$. During the exploration process, as the map is updated, the $\mathbf{SI_{i}}$ of $S_{i}$ also changes accordingly. Therefore, all cells intersecting with the map update region $\mathcal{B}_{u}$ need to recompute their SIS. Algorithm 1 outlines the online subregion cell updating process. First, traverse the set $\mathcal{V}$ to filter the cell $S_{i}$ that overlaps with $\mathcal{B}_{u}$ (Line 2). Then, some information such as the proportion of known voxels $\mathbf{R}_{\text{known},i}$ and the status $stat_{i}$ is updated (Line 4). For $S_{i}$ in active status, if $\mathbf{R}_{\text{known},i}$ exceeds a certain threshold $\sigma_{k}$ and is not at $l_{\text{max}}$, it is subdivided into four subcells (Line 8). Then, the $S_{i}$ is removed from $\mathcal{V}$, and the SIS of the subcells is calculated (Lines 9-12). Otherwise, the passable path from $S_{i}$ to other subregion cells is searched in $\mathcal{G}$ to update traveling cost $\mathbf{L}_{\text{cost},i}$ (Line 14). Note that the update of $\mathbf{L}_{\text{cost}}$ between subregion pair is mutual, so the traveling cost from another subregion $S_{j}$ to $S_{i}$ also needs to be updated, and $\mathbf{L}_{\text{cost},i,j}=\mathbf{L}_{\text{cost},j,i}$.
For $N_{\text{act}}$ subregions in active status, if we calculate the traveling cost $\mathbf{L}_{\text{cost}}$ between all subregion pairs, $O(N_{\text{act}}^{2})$ A* path searches are required at each iteration. In most cases, the number of $N_{\text{act}}$ is not large. However, in extensive environments, this can still be very computationally intensive and may cause the robot to freeze if the subregion pairs are far apart. Theoretically, only the subregion cells intersecting the updated region $\mathcal{B}_{u}$ are affected, so only the cost data of these cells need to be updated. Assume that $M$ subregion cells are affected, in this case, only $O(MN_{\text{act}})$ path searches are performed. Practically, the number of $M$ is small, resulting in a small number of path search computations.
// Initialize $\mathcal{V}=\left\{S_{o}\right\}$ at the beginning
@@
2       if $\neg$ isOverlap($S_{i},\mathcal{B}_{u}$) then
3             continue
4       end if
5      updateInfo($S_{i}$)
6       if $\neg$ isActive($S_{i}$) then
7             continue
8       end if
9      if $\mathbf{R}_{\text{known},i}>\sigma_{k}\land l_{i}<l_{\text{max}}$ then
10             $S_{\text{div}}\leftarrow$ subdivideCell($S_{i}$)
11             $\mathcal{V}$.erase($S_{i}$)
12             for $S_{l}\in S_{\text{div}}$ do
13                   calculateSIS($S_{l}$)
14                   $\mathcal{V}$.append($S_{l}$)
15
@@ VII-B Ablation Experiment
To study the specific impact of the greedy strategy on exploration efficiency, we replace the hierarchical planning strategy with the mechanism of moving to the nearest viewpoint as proposed in [6]. The results are shown in Table. III. During exploration, the greedy strategy leads to frequent back-and-forth movements between several viewpoints, reducing exploration efficiency. Furthermore, this strategy can result in some areas being missed, necessitating additional retracing to ensure complete coverage. Therefore, it is crucial to assess the travel cost of visiting all areas to be explored. This allows for arranging an efficient sequence for visiting different regions, providing overall guidance for the entire exploration process.
We also investigate the impact of distance calculation methods on solving the TSP. We use Euclidean distance to describe the distance between two areas instead of the exact distance obtained by A* search. In constructing the TSP distance matrix, we find that inaccurate distance measurements not only lead to suboptimal TSP solutions but also cause frequent path switching, thereby affecting overall optimization performance. Therefore, choosing an appropriate distance measurement is crucial for effectively solving the TSP. Accurate distance measurements can precisely reflect travel costs, ensuring a reasonable visitation sequence.
TABLE III: Evaluation of Two Mechanisms on Exploration Performance
Scene Method Distance(m) Time Duration(s)
Avg Std Avg Std
Scene 1 Greedy Strategy 1167.10 15.12 763.30 32.20
Euclidean Distance 1419.75 12.69 998.09 24.00
Scene 2 Greedy Strategy 1548.78 34.6 865.78 20.65
Euclidean Distance 1606.65 27.5 949.68 24.98

### v3

@@ Abstract
We present an autonomous exploration system for efficient coverage of unknown environments. First, a rapid environment preprocessing method is introduced to provide environmental information for subsequent exploration planning. The exploration space is then divided into multiple subregion cells with varying levels of detail, capable of online decomposition and updating to adaptively represent dynamic unknown regions. Finally, a hierarchical planning strategy treats subregions as basic units, computing an efficient global coverage path. The local path is refined under the guidance of the global path to sequentially visit the viewpoint set, providing an executable trajectory for the robot. This coarse-to-fine planning reduces complexity while enhancing exploration efficiency. Compared with state-of-the-art methods in benchmark environments, the proposed approach achieves superior exploration efficiency with lower computational costs.
@@ I Introduction
Autonomous exploration using ground mobile robots has received considerable attention in recent years for its wide applications, including searching and rescuing, environmental monitoring, and other tasks. The ability of robots to navigate and explore unknown environments without human intervention enhances operational efficiency and ensures personnel safety in hazardous conditions. Although several autonomous exploration methods have been proposed over the decades [1, 2, 3], the task remains challenging in large-scale and topologically complex scenarios. In such scenarios, the problem becomes computationally complex, and the robot needs to navigate between areas in a limited time while avoiding redundant exploration.
The mainstream exploration frameworks mainly include frontier-based and sampling-based methods. In autonomous exploration, the frontier is defined as the boundary between free and unknown areas [1]. These two approaches guide the robot to cover the environment by continuously searching for frontiers in the map or randomly sampling viewpoints and evaluating their potential revenue. However, several limitations still hinder the development of autonomous exploration:
@@
Figure 1: The real-world experiment conducted in an underground garage using real ground platforms. The yellow dot denotes the start position, and the white line indicates the robot’s trajectory. Video of the experiment is available at: https://youtu.be/5shy7c1Faf4.
In our previous work [9], the hierarchical planning strategy demonstrates significant advantages in autonomous exploration. However, during space division, all subregions use the same resolution, which limits their ability to adapt to variations in the scale and structure of unknown areas, thereby reducing the capability for refined representation. Furthermore, the local planning process relies on a reward-based strategy, which results in shortsightedness, reduced efficiency in large-scale environments, and challenges in ensuring global optimization in complex scenarios.
In this work, we propose an efficient hierarchical exploration planner for autonomous exploration in complex and large-scale environments. First, an incremental environmental preprocessing method is introduced to generate a set of high-quality viewpoints as candidate target points. To better capture the dynamics of unknown areas during exploration, a variable-resolution regional division approach is applied, enabling subregion cells to be further subdivided and adjusted as the unknown areas are transformed into known areas during exploration. Additionally, the Subregion Information Structure (SIS) is designed to record essential information within the subregion space, facilitating exploration planning. Finally, subregions are employed as the basic planning units, and a hierarchical planning strategy is developed for efficient exploration. The global coverage path determines the shortest travel route that sequentially visits each subregion, while the local path refines viewpoint access under the guidance of the global path. This process is continuously executed until the entire environment is fully explored.
@@
  • A subregion-based regional division method with a hierarchical structure is proposed, enabling adaptive decomposition of subregions to effectively represent the evolving characteristics of unknown regions.
@@ II Related Works
Several methods [14], [15], [16], [17] are proposed for the concepts of regional division and hierarchical planning. However, these methods are either designed for task allocation in multi-robot collaborative exploration or utilize fixed-sized regions that cannot adapt to the scale and structural changes of the unknown region. In [16], the global path is introduced to provide a more comprehensive perspective for exploration. However, the greedy strategy is still adopted in the local planning stage, which can still lead to short-sighted behavior and thus reduce exploration efficiency. The concept of region hierarchical division in this paper is inspired by [15], which utilizes hgrid space for muti-UAV interaction and task assignment. Our approach of the regional division is aimed at serving the global coverage path planning, which is part of the hierarchical planning.
Learning-based techniques have also emerged as promising solutions for autonomous exploration [18], [19]. These methods leverage data-driven models to predict optimal exploration strategies but often lack robustness and adaptability to unfamiliar scenarios. This leads to performance deterioration in complex and unfamiliar environments.
@@ IV-A Mapping and Frontier Detection
As in state-of-the-art works [11, 12], an incremental frontier detection method is employed in this work to avoid processing the entire map. Specifically, a set $\mathcal{C}_{v}$ of voxels with updated occupation values is obtained from the UFOMap, and the state of each voxel $V\in\mathcal{C}_{v}$ is checked. If a voxel $V$ is free ($V_{free}$) and has an adjacent voxel in the unknown state ($V_{un}$) at the same height, it is identified as a frontier voxel. During exploration, each frontier voxel in the map update region $\mathcal{B}_{u}$ is checked to verify its validity, and invalid frontiers are removed. This method requires examining only a small subset of voxels, significantly reducing the workload for frontier extraction.
@@ IV-B Dynamic Sparse Graph
Given that path search operations are time-consuming in the 3D voxel map, a dynamic sparse graph $\mathcal{G}=(N,E)$ is maintained as a sparse environment representation, where $N$ denotes nodes and $E$ represents undirected weighted edges. Candidate nodes $N_{c}$ are generated via biased random sampling based on the terrain map $\mathcal{A}$, and valid nodes are added to $\mathcal{G}$ after collision checking. Edge weights are set as the Euclidean distance between nodes, and all nodes are stored in an R*-tree[21] to enhance computational efficiency.
@@
where ${Accessible}(\mathcal{A})$ represents the accessible region in $\mathcal{A}$, ensuring all nodes are reachable by the robot. The parameters $\delta$ and $\gamma$ define the minimum and maximum allowable distances between $N_{c}$ and the nearest node $N_{\text{nearest}}$ in $\mathcal{G}$, preventing the graph from becoming overly dense or sparse while maintaining shortest paths between nodes.
Since previously passable paths can be blocked by dynamic obstacles such as falling rocks, failing to update the sparse graph may lead to incorrect paths or inaccurate traversal distances, negatively impacting exploration efficiency and safety. To address this, the sparse graph is incrementally maintained within the updated region $\mathcal{B}_{u}$. Specifically, nodes and edges within $\mathcal{B}_{u}$ are subject to collision detection, with invalid edges caused by environmental changes being removed and valid nodes and edges being added. This method improves graph update efficiency and significantly reduces the computational cost of global graph rebuilding, enabling more efficient and robust exploration in dynamic environments.
@@ IV-C Viewpoint Generation
where $\mathbf{p}_{o}$ is the position of the nearest occupied voxel, and $r_{\text{sensor}}$ is the maximum scanning radius of the LiDAR sensor. Due to the absence of obstacles within the sphere and its convex nature, any frontier voxel in the sphere can be observed from a viewpoint inside. Thus, performing expensive ray tracing on every frontier voxel to determine its visibility from the viewpoint is unnecessary. Then, the node $N$ located inside the sphere and closest to its center is selected as the viewpoint. This search process is achieved by K-Nearest Neighbor search in $\mathcal{G}$. Fig. 3 shows the dynamic sparse graph and the viewpoints generated based on collision-free spheres.
@@ V Subregion-based Regional Division
Instead of directly determining the visit order of viewpoints or frontiers, we employ subregion cells as planning units for global coverage. Each cell, managed through online decomposition and updates, enhances its ability to adapt to the scale and structural changes of unknown regions with multi-resolution.
@@ V-B Subregion Information Structure
The Subregion Information Structure $\mathbf{SI_{i}}$ is composed of elemental information contained in each subregion cell $S_{i}$. The hierarchy level $l_{i}$ and axis-aligned bounding box $\mathbf{B}_{i}$ are recorded when initializing $S_{i}$. The proportion of known voxels $\mathbf{R}_{\text{known},i}$ and the viewpoint average position $\mathbf{C}_{\text{avg},i}$ of $S_{i}$ are also computed to locate the area to be explored. To serve subsequent global coverage path planning, the traveling cost $\mathbf{L}_{\text{cost},i}$ between $S_{i}$ and other subregions is also computed. The viewpoints $\mathbf{VP}_{i}$ within $S_{i}$ are stored in $\mathbf{SI_{i}}$ for subsequent local path optimization. In addition, each subregion cell holds a status $stat_{i}$ between active and inactive. Subregion cells containing at least one viewpoint are categorized as active, while those without viewpoints are considered inactive. Only subregions with active status are eligible for inclusion in the global coverage path planning process. Table. I presents the meaning of each data information in SIS.
@@
$stat_{i}$ active or inactive status of $S_{i}$
@@ V-C Online Subregion Updating
We maintain a set $\mathcal{V}$ that stores all divided subregions. Initially, the entire unknown space $\mathcal{O}$ is divided into $S_{o}=S_{o,x}\times S_{o,y}$ cells at level $l=1$. During exploration, as the map updates, the subregion information $\mathbf{SI_{i}}$ of $S_{i}$ is modified accordingly. Therefore, all subregion cells intersecting with the map update region $\mathcal{B}_{u}$ need to recompute their SIS. Algorithm 1 outlines the online subregion cell updating process. First, traverse the set $\mathcal{V}$ to filter the cell $S_{i}$ that overlaps with $\mathcal{B}_{u}$ (Line 3). Then, some attributes such as the proportion of known voxels $\mathbf{R}_{\text{known},i}$ and the status $stat_{i}$ is updated (Line 4). For $S_{i}$ in active status, if $\mathbf{R}_{\text{known},i}$ exceeds a threshold $\sigma_{k}$ and $l_{i}<l_{\text{max}}$, the cell is subdivided, removed from $\mathcal{V}$, and its subcells are added after recalculating their SIS (Lines 5–12). Otherwise, passable paths from $S_{i}$ to other subregions are updated using the graph search in $\mathcal{G}$ (Line 15). Since the traveling cost $\mathbf{L}_{\text{cost},i,j}$ is symmetric, it is updated for both $i\to j$ and $j\to i$.
1 Initialize $\mathcal{V}=\{S_{o}\}$ ;
@@
3       if overlapCheck($S_{i},\mathcal{B}_{u}$) then
4             updateAttributes($S_{i}$) ;
5             if isActive($S_{i}$) then
6                   if $\mathbf{R}_{\text{known},i}>\sigma_{k}$ and $l_{i}<l_{\text{max}}$ then
7                         $S_{\text{new}}\leftarrow\texttt{splitRegion}(S_{i})$ ;
8                         $\mathcal{V}\leftarrow\mathcal{V}\setminus\{S_{i}\}$ ;
9                         for $S_{j}\in S_{\text{new}}$ do
10                               computeSIS($S_{j}$) ;
11                               $\mathcal{V}\leftarrow\mathcal{V}\cup\{S_{j}\}$ ;
12
@@ VII-B Ablation Experiment
As outlined in Sec.V-A, a variable-resolution mechanism is employed to enhance adaptability to the scale and structural changes of the unknown region. However, how do different resolution settings specifically affect exploration efficiency? Moreover, to what extent do these settings influence the computational requirements of the proposed approach?
We design an experimental variation using two fixed resolutions ($W=8m$ and $W=25m$) for subregion division. The experimental results, as shown in Table. III, indicate that subregions with finer resolution achieve better exploration performance. This result aligns with intuition, as higher-resolution representations provide more precise global coverage paths. However, this also significantly increases the complexity of solving the TSP, leading to higher computation times, which deviates from the intended efficiency of our coarse-to-fine planning framework. Conversely, overly coarse resolutions fail to provide effective global exploration guidance, thereby slowing down the overall exploration process.
In contrast, the hierarchical subregion structure dynamically adjusts the spatial size of subregions based on environmental characteristics, enabling flexible representation of unknown areas. Finer subdivisions are applied to large covered regions to enhance exploration precision, while lower-resolution representations are employed for vast uncovered regions to reduce computational overhead. This dynamic adjustment strategy not only accurately captures the features of unexplored areas but also significantly reduces the total number of subregions, achieving a desirable balance between exploration efficiency and computational cost.
TABLE III: Impact of Subregion Resolutions on Exploration Performance
Scene Resolution Distance(m) Time Duration(s) Run Time(s)
Avg Std Avg Std Avg Std
Scene 1 ($100\times 80$) $W=8m$ 974.34 14.37 587.27 8.87 0.245 0.124
$W=25m$ 1121.70 15.32 655.88 10.41 0.096 0.068
$Variable$ 992.62 10.13 589.60 5.25 0.125 0.094
Scene 2 ($137\times 104$) $W=8m$ 1440.11 30.45 818.96 21.77 0.182 0.098
$W=25m$ 1597.68 26.98 915.49 16.41 0.063 0.045
$Variable$ 1439.82 29.39 819.56 13.22 0.085 0.072
Scene 3 ($80\times 80$) $W=8m$ 611.85 25.18 377.13 18.87 0.122 0.089
$W=25m$ 756.67 19.76 478.78 23.21 0.068 0.037
$Variable$ 630.76 27.13 390.40 20.10 0.075 0.043